# TimeBench: A Comprehensive Evaluation of Temporal Reasoning Abilities in Large Language Models

### v1

@@ A.4 Prompts
Context: {}
Question: {}
Options: {}
Answer:
TimeDial
There is a two-person dialogue with several options.
Choose all appropriate options to substitute the <mask> in the dialogue, and each question has at least one correct option.
Dialogue: {}
Options: {}
Answer:
TRACIE
Read the following story and hypothesis, determine whether the hypothesis can be inferred from the story.
You need to understand the implicit temporal relationships between events to make judgments.
Story: {}
Hypothesis: {}
Options: A. Entailment B. Contradiction
Answer:
SituatedGen
Generate a pair of contrastive sentences with the given set of keywords.
Keywords: {}
Date Arithmetic
Question: {}? Answer:
TimeQA
I will give you a question with context.
You need to answer my question based on the context.
If you can infer the answer from the context, then output your answer. Otherwise, if there is no answer, output [unanswerable].
Context: {}
Question: {}
Answer:
TempReason
I will give you a question with context.
You need to answer my question based on the context.
Context: {}
Question: {}
Answer:
MenatQA
Get answers for the question based on the contxt, where answers derived from substrings in the context or categorized as [unanswerable].
Context: {}
Question: {}
Answer:
TimeX-NLI
Read the following statements about time and determine if the hypothesis can be inferred from the premise.
Premise: {}
Hypothesis: {}
Options: A. Entailment B. Contradiction C. Neutral
Answer:
Figure 6: Zeroshot instructions and input formats.
@@
Premise: On Wednesday, they got married.
Hypothesis: Before Friday, they got married.
Options: A. Entailment B. Contradiction C. Neutral
Answer: Wednesday is before Friday. As a result, we can infer that if something happens on Wednesday, it definitely happens before Friday. Therefore, the answer is A. Entailment.
Premise: We went to Disneyland on Monday.
Hypothesis: We went to Disneyland after Wednesday.
Options: A. Entailment B. Contradiction C. Neutral
Answer: Monday is before Wednesday. As a result, We can infer that if something happens on Monday, it definitely can not happen after Wednesday. Therefore, the answer is B. Contradiction.
Premise: The failing company issued major layoffs after Tuesday.
Hypothesis: The failing company issued major layoffs after Thursday.
Options: A. Entailment B. Contradiction C. Neutral
Answer: Tuesday is before Thursday. If something happened after Tuesday, we cannot be certain whether it occurred after Thursday. Therefore, the answer is C. Neutral.
CoT Demonstration of Date Arithmetic (4-shot)
Question: What is the time 4 year and 1 month after Apr, 2000?
Answer: First, 4 years after 2000 is 2004. Next, 1 month after April is May. Therefore, 4 year and 1 month after Apr, 2000 is May, 2004.
Question: What is the time 3 year and 4 month before Jun, 1840?
Answer: First, subtracting 3 years from 1840 gives 1837. Next, subtracting 4 months from June gives February. Therefore, 3 year and 4 month before Jun, 1840 is Feb, 1837.
Question: What is the time 7 year and 11 month after Feb, 1819?
Answer: First, 7 years after 1819 is 1826. Next, 11 months after February is January of the next year. Therefore, 7 years and 11 months after Feb, 1819 is Jan, 1827.
Question: What is the time 6 year and 9 month before Jan, 1234?
Answer: First, subtracting 6 years from 1234 gives 1228. Next, subtracting 9 months from January gives April of the previous year. Therefore, 6 year and 9 month before Jan, 1234 is Apr, 1227.
Figure 8: Chain-of-Thought demonstrations of Date Arithmetic.
CoT Demonstration of TRACIE (4-shot)
Read the following story and hypothesis, determine whether the hypothesis can be inferred from the story.
You need to understand the implicit temporal relationships between events to make judgments
Story: Joe was a police officer. Joe was patrolling the streets of the city in his cruiser. S̈uddenly, Joe was alerted to a crime happening near him by dispatch.J̈oe responded to the scene and found a bank robber fleeing on foot. Joe arrested the criminal and was promoted.
Hypothesis: Joe put on his police uniform. starts after Joe arrest the criminal
Options: A. Entailment B. Contradiction
Answer: From the story we know Joe was patrolling. In the work state, Joe has already put on the police uniform. So we can infer that Joe put on his police uniform before arresting the criminal. This conflicts with hypothesis. Therefore, the answer is B. Contradiction.
Figure 9: Chain-of-Thought demonstrations of TRACIE.
@@
Context: actually i have an project on it so please give me as much as you have information about migratory birds in punjab
Question: How long did it take for them to have information about migratory birds in punjab?
Options: A. several months B. 12 weeks C. a few minutes D. almost instantly
Answer: This is a conversation scenario. In the conversation, providing relevant information about migratory birds in punjab to him is in real-time and takes very little time. Therefore, the answer is C. a few minutes, D. almost instantly.
Context: Hope she stops laying eggs because she will get really skinny !
Question: How long did it take for her to lay eggs?
Options: A. 1 week B. 22 hours C. 2 years D. 4 years
Answer: According to commonsense knowledge, the time it takes for birds to lay eggs typically varies from one day to several days. Therefore, the answer is A. 1 week, B. 22 hours.
Figure 10: Chain-of-Thought demonstrations of DurationQA.
@@
Context: She ordered the tastiest kind of each vegetable and the prettiest kind of each flower.
Question: How often does she order vegetables and flowers?
Options: A. once a second B. three days a week C. every 10 centuries D. once a week
Answer: According to commonsense knowledge, ordering vegetables and flowers typically happens on a regular basis, usually every few days. Therefore, the answer is B. three days a week, D. once a week.
Context: Wallace, 38, called Gastonia home from the age of 8 until she graduated from Hunter Huss High School in 1983.
Question: When did Wallace wake up for high school?
Options: A. at 6 am B. at 1 am C. 7:00 AM D. at 6 pm
Answer: According to commonsense knowledge, waking up for high school typically happens in the morning, usually between 6 AM and 8 AM. Therefore, the answer is A. at 6 am, C. 7:00 AM.
@@
There is a two-person dialogue with several options.
Choose all appropriate options to substitute the <mask> in the dialogue, and each question has at least one correct option.
Dialogue:
A:What schools have you attended ?
B: I finished Young Primary School in 1998 , and entered Xi ’ an Middle School that same September . I graduated from there in <MASK> , and that September I entered Wuhan University , where I’m studying now .
A: How do you think the education you have received will contribute to your work in this company ?
B: I think I have a good understanding of fundamentals in the areas your company deals with , and I can go on from here to build up the specific skills and knowledge I need to do my job well .
A: Your graduation thesis was on Medical Application of Laser , right ? What were your conclusions ?
B: Yes . I did some work on that , and I found out some really interesting things about the conductivity of liquid helium . I was sure I had a great discovery until my teacher told me the same discovery already made twenty years ago . I think the most important thing , I learnt though , was the importance of keeping good records .
Options: A. 1998 B. July of 2004 C. March of 2003 D. twenty years ago
Answer: Based on the dialogue, B entered middle school in Sep 1998. According to commonsense knowledge, it usually takes around 6 years from entering middle school to graduating from high school (and entering university). Adding 6 years to 1998 would be 2004, so the answer should be around the year 2004. Therefore, the answer is B. July of 2004, C. March of 2003.
@@
I will give you a question with context.
You need to answer my question based on the context.
If you can infer the answer from the context, then output your answer. Otherwise, if there is no answer, output [unanswerable]
Context: Theo-Ben Gurirab Theo-Ben Gurirab ( 23 January 1938 2̆013 14 July 2018 ) was a Namibian politician who served in various senior government positions . He served as the second Prime Minister of Namibia from 28 August 2002 to 20 March 2005 , following the demotion and subsequent resignation of Hage Geingob . Previously he was the countrys first Minister of Foreign Affairs from 1990 to 2002 , and was President of the United Nations General Assembly from 1999 to 2000 . He was Speaker of the National Assembly of Namibia from 2005 to 2015 , when he was replaced by Peter Katjavivi . Gurirab ultimately resigned from politics in 2015 . Death . Gurirab died at a Windhoek hospital on 14 July 2018 of natural causes . He is buried at Heroes Acre .
Question: Theo-Ben Gurirab took which position after Jan 2007?
Answer: Based on the context, we can summarize the following facts: Theo-Ben Gurirab served as second Prime Minister of Namibia from August 2002 to March 2005. Prior to that, he was the countrys first Minister of Foreign Affairs from 1990 to 2002 and and was President of the United Nations General Assembly from 1999 to 2000. From 2005 to 2015, he held the position of Speaker of the National Assembly of Namibia. He resigned from politics in 2015 and passed away in July 2018. According to the aforementioned facts, he took the position of Speaker of the National Assembly of Namibia in January 2007. Therefore, the answer is Speaker of the National Assembly of Namibia.
@@
I will give you a question with context.
You need to answer my question based on the context.
Context (facts): Gian Piero Gasperini is the head coach of Atalanta B.C. from Jun, 2016 to Dec, 2022.
Edoardo Reja is the head coach of Atalanta B.C. from Mar, 2015 to Jun, 2016.
Stefano Colantuono is the head coach of Atalanta B.C. from Jun, 2010 to Mar, 2015.
Bortolo Mutti is the head coach of Atalanta B.C. from Jan, 2010 to Jun, 2010.
Emiliano Mondonico is the head coach of Atalanta B.C. from Jul, 1987 to Jun, 1990.
Marcello Lippi is the head coach of Atalanta B.C. from Jul, 1992 to Jun, 1993.
Angelo Gregucci is the head coach of Atalanta B.C. from Jul, 2009 to Sep, 2009.
Luigi Delneri is the head coach of Atalanta B.C. from Jul, 2007 to Jun, 2009.
Ottavio Bianchi is the head coach of Atalanta B.C. from Jul, 1981 to Jun, 1983.
Antonio Conte is the head coach of Atalanta B.C. from Sep, 2009 to Jan, 2010.
Nedo Sonetti is the head coach of Atalanta B.C. from Jul, 1983 to Jun, 1987.
Valter Bonacina is the head coach of Atalanta B.C. from Jan, 2010 to Jan, 2010.
Question: Who was the head coach of the team Atalanta B.C. in Feb, 2016?
Answer: According to the context, Edoardo Reja was the head coach of Atalanta B.C. from Mar, 2015 to Jun, 2016. In Feb 2016, the head coach of the team Atalanta B.C. is Edoardo Reja. Therefore, the answer is Edoardo Reja.
@@
I will give you a question with context.
You need to answer my question based on the context.
Context (facts): Nicholas Macpherson holds the position of Member of the House of Lords from Oct, 2016 to Dec, 2022.
Nicholas Macpherson holds the position of Principal Private Secretary to the Chancellor of the Exchequer from Jan, 1993 to Jan, 1997.
Nicholas Macpherson holds the position of Permanent Secretary to the Treasury from Aug, 2005 to Jan, 2016.
Question: Which position did Nicholas Macpherson hold before Member of the House of Lords?
Answer: According to the context, Nicholas Macpherson holds the position of Permanent Secretary to the Treasury from Aug, 2005 to Jan, 2016. Afterthat, Nicholas Macpherson holds the position of Member of the House of Lords from Oct, 2016 to Dec, 2022. Nicholas Macpherson hold the position of Permanent Secretary to the Treasury before Member of the House of Lords. Therefore, the answer is Permanent Secretary to the Treasury."

### v2

@@ B.3 Prompts
Premise: On Wednesday, they got married. Hypothesis: Before Friday, they got married. Options: A. Entailment B. Contradiction C. Neutral Answer: Wednesday is before Friday. As a result, we can infer that if something happens on Wednesday, it definitely happens before Friday. Therefore, the answer is A. Entailment.
Premise: We went to Disneyland on Monday. Hypothesis: We went to Disneyland after Wednesday. Options: A. Entailment B. Contradiction C. Neutral Answer: Monday is before Wednesday. As a result, We can infer that if something happens on Monday, it definitely can not happen after Wednesday. Therefore, the answer is B. Contradiction.
Premise: The failing company issued major layoffs after Tuesday. Hypothesis: The failing company issued major layoffs after Thursday. Options: A. Entailment B. Contradiction C. Neutral Answer: Tuesday is before Thursday. If something happened after Tuesday, we cannot be certain whether it occurred after Thursday. Therefore, the answer is C. Neutral.
@@
Question: What is the time 4 year and 1 month after Apr, 2000? Answer: First, 4 years after 2000 is 2004. Next, 1 month after April is May. Therefore, 4 year and 1 month after Apr, 2000 is May, 2004.
Question: What is the time 3 year and 4 month before Jun, 1840? Answer: First, subtracting 3 years from 1840 gives 1837. Next, subtracting 4 months from June gives February. Therefore, 3 year and 4 month before Jun, 1840 is Feb, 1837.
Question: What is the time 7 year and 11 month after Feb, 1819? Answer: First, 7 years after 1819 is 1826. Next, 11 months after February is January of the next year. Therefore, 7 years and 11 months after Feb, 1819 is Jan, 1827.
Question: What is the time 6 year and 9 month before Jan, 1234? Answer: First, subtracting 6 years from 1234 gives 1228. Next, subtracting 9 months from January gives April of the previous year. Therefore, 6 year and 9 month before Jan, 1234 is Apr, 1227.
CoT Demonstration of Date Arithmetic (4-shot)
Figure 11: Chain-of-Thought demonstrations of Date Arithmetic.
Read the following story and hypothesis, determine whether the hypothesis can be inferred from the story. You need to understand the implicit temporal relationships between events to make judgments
Story: Joe was a police officer. Joe was patrolling the streets of the city in his cruiser. S̈uddenly, Joe was alerted to a crime happening near him by dispatch.J̈oe responded to the scene and found a bank robber fleeing on foot. Joe arrested the criminal and was promoted. Hypothesis: Joe put on his police uniform. starts after Joe arrest the criminal Options: A. Entailment B. Contradiction Answer: From the story we know Joe was patrolling. In the work state, Joe has already put on the police uniform. So we can infer that Joe put on his police uniform before arresting the criminal. This conflicts with hypothesis. Therefore, the answer is B. Contradiction.
CoT Demonstration of TRACIE (4-shot)
Figure 12: Chain-of-Thought demonstrations of TRACIE.
@@
Context: actually i have an project on it so please give me as much as you have information about migratory birds in punjab Question: How long did it take for them to have information about migratory birds in punjab? Options: A. several months B. 12 weeks C. a few minutes D. almost instantly Answer: This is a conversation scenario. In the conversation, providing relevant information about migratory birds in punjab to him is in real-time and takes very little time. Therefore, the answer is C. a few minutes, D. almost instantly.
Context: Hope she stops laying eggs because she will get really skinny ! Question: How long did it take for her to lay eggs? Options: A. 1 week B. 22 hours C. 2 years D. 4 years Answer: According to commonsense knowledge, the time it takes for birds to lay eggs typically varies from one day to several days. Therefore, the answer is A. 1 week, B. 22 hours.
@@
Context: She ordered the tastiest kind of each vegetable and the prettiest kind of each flower. Question: How often does she order vegetables and flowers? Options: A. once a second B. three days a week C. every 10 centuries D. once a week Answer: According to commonsense knowledge, ordering vegetables and flowers typically happens on a regular basis, usually every few days. Therefore, the answer is B. three days a week, D. once a week.
Context: Wallace, 38, called Gastonia home from the age of 8 until she graduated from Hunter Huss High School in 1983. Question: When did Wallace wake up for high school? Options: A. at 6 am B. at 1 am C. 7:00 AM D. at 6 pm Answer: According to commonsense knowledge, waking up for high school typically happens in the morning, usually between 6 AM and 8 AM. Therefore, the answer is A. at 6 am, C. 7:00 AM.
@@
There is a two-person dialogue with several options. Choose all appropriate options to substitute the <mask> in the dialogue, and each question has at least one correct option.
Dialogue: A:What schools have you attended ? B: I finished Young Primary School in 1998 , and entered Xi ’ an Middle School that same September . I graduated from there in <MASK> , and that September I entered Wuhan University , where I’m studying now . A: How do you think the education you have received will contribute to your work in this company ? B: I think I have a good understanding of fundamentals in the areas your company deals with , and I can go on from here to build up the specific skills and knowledge I need to do my job well . A: Your graduation thesis was on Medical Application of Laser , right ? What were your conclusions ? B: Yes . I did some work on that , and I found out some really interesting things about the conductivity of liquid helium . I was sure I had a great discovery until my teacher told me the same discovery already made twenty years ago . I think the most important thing , I learnt though , was the importance of keeping good records . Options: A. 1998 B. July of 2004 C. March of 2003 D. twenty years ago Answer: Based on the dialogue, B entered middle school in Sep 1998. According to commonsense knowledge, it usually takes around 6 years from entering middle school to graduating from high school (and entering university). Adding 6 years to 1998 would be 2004, so the answer should be around the year 2004. Therefore, the answer is B. July of 2004, C. March of 2003.
@@
I will give you a question with context. You need to answer my question based on the context. If you can infer the answer from the context, then output your answer. Otherwise, if there is no answer, output [unanswerable]
Context: Theo-Ben Gurirab Theo-Ben Gurirab ( 23 January 1938 2̆013 14 July 2018 ) was a Namibian politician who served in various senior government positions . He served as the second Prime Minister of Namibia from 28 August 2002 to 20 March 2005 , following the demotion and subsequent resignation of Hage Geingob . Previously he was the countrys first Minister of Foreign Affairs from 1990 to 2002 , and was President of the United Nations General Assembly from 1999 to 2000 . He was Speaker of the National Assembly of Namibia from 2005 to 2015 , when he was replaced by Peter Katjavivi . Gurirab ultimately resigned from politics in 2015 . Death . Gurirab died at a Windhoek hospital on 14 July 2018 of natural causes . He is buried at Heroes Acre . Question: Theo-Ben Gurirab took which position after Jan 2007? Answer: Based on the context, we can summarize the following facts: Theo-Ben Gurirab served as second Prime Minister of Namibia from August 2002 to March 2005. Prior to that, he was the countrys first Minister of Foreign Affairs from 1990 to 2002 and and was President of the United Nations General Assembly from 1999 to 2000. From 2005 to 2015, he held the position of Speaker of the National Assembly of Namibia. He resigned from politics in 2015 and passed away in July 2018. According to the aforementioned facts, he took the position of Speaker of the National Assembly of Namibia in January 2007. Therefore, the answer is Speaker of the National Assembly of Namibia.
@@
I will give you a question with context. You need to answer my question based on the context.
Context (facts): Gian Piero Gasperini is the head coach of Atalanta B.C. from Jun, 2016 to Dec, 2022. Edoardo Reja is the head coach of Atalanta B.C. from Mar, 2015 to Jun, 2016. Stefano Colantuono is the head coach of Atalanta B.C. from Jun, 2010 to Mar, 2015. Bortolo Mutti is the head coach of Atalanta B.C. from Jan, 2010 to Jun, 2010. Emiliano Mondonico is the head coach of Atalanta B.C. from Jul, 1987 to Jun, 1990. Marcello Lippi is the head coach of Atalanta B.C. from Jul, 1992 to Jun, 1993. Angelo Gregucci is the head coach of Atalanta B.C. from Jul, 2009 to Sep, 2009. Luigi Delneri is the head coach of Atalanta B.C. from Jul, 2007 to Jun, 2009. Ottavio Bianchi is the head coach of Atalanta B.C. from Jul, 1981 to Jun, 1983. Antonio Conte is the head coach of Atalanta B.C. from Sep, 2009 to Jan, 2010. Nedo Sonetti is the head coach of Atalanta B.C. from Jul, 1983 to Jun, 1987. Valter Bonacina is the head coach of Atalanta B.C. from Jan, 2010 to Jan, 2010. Question: Who was the head coach of the team Atalanta B.C. in Feb, 2016? Answer: According to the context, Edoardo Reja was the head coach of Atalanta B.C. from Mar, 2015 to Jun, 2016. In Feb 2016, the head coach of the team Atalanta B.C. is Edoardo Reja. Therefore, the answer is Edoardo Reja.
@@
I will give you a question with context. You need to answer my question based on the context.
Context (facts): Nicholas Macpherson holds the position of Member of the House of Lords from Oct, 2016 to Dec, 2022. Nicholas Macpherson holds the position of Principal Private Secretary to the Chancellor of the Exchequer from Jan, 1993 to Jan, 1997. Nicholas Macpherson holds the position of Permanent Secretary to the Treasury from Aug, 2005 to Jan, 2016. Question: Which position did Nicholas Macpherson hold before Member of the House of Lords? Answer: According to the context, Nicholas Macpherson holds the position of Permanent Secretary to the Treasury from Aug, 2005 to Jan, 2016. Afterthat, Nicholas Macpherson holds the position of Member of the House of Lords from Oct, 2016 to Dec, 2022. Nicholas Macpherson hold the position of Permanent Secretary to the Treasury before Member of the House of Lords. Therefore, the answer is Permanent Secretary to the Treasury."